PIC1 - Completed OK after a the solution of an adjustment problem in the current loop controller of the QPS (few mA too low).
Connection of the 60A conveter to the RB circuit in the UA83 side. The TF measurements could not be completed because the converter noise made the bus-bar quench detector trip everytime it was switched on. The detection threshold will be increased on Monday.
The time out preventing from closing the switches too early after a switch opening has been lowered in order to speed the tests up.
RQF.A78
PLI1 (760A) completed - Test OK
PLI2 (2000A) completed - Test OK
PLI3.1 (current cycle up to 6500A) completed - Test OK
60 Closed Orbit Correctors
PCC carried out in one circuit.
All the circuits released by ElQA have been connected to the power converter.
AB/PO current stability measurements in RD2.L8 - Completed
Empty of DFBMA and Q4/D2 started in the evening
Test Plan for Monday 25.06.2007
RB.A78
Provisional correction of the detection threshold of the bus-bar quench detector in order to allow TF measurement with the 60A converter.
TF measurement
Desconnection of the 60A converter and final connection of the RB converter to the circuit; PCC tests
Set detection threshold back to operation value
PIC2
PLI1 (760A)
PLI1.1 - Current cycle
PLI1.2 - Energy extraction discharge (quench to a current lead)
PLI1.3 - Powering Failure from PC
PLI1.4 - FPA from PIC
PLI1.5 - Quench heater discharge on a selected magnet
PLI2 (2000A)
PLI2.1 - Current cycle
PLI2.2 - Energy extraction discharge (quench to a current lead)
PLI2.2b - SPA from PIC
PLI2.3 - Quench heater discharge on a selected magnet
RQF.A78 (6500A)
PLI3.2 - Energy extraction discharge
PLI3.3 - Quench heater discharge on 12L8 SSS
600A Correctors
Software upgrade by QPS
PCS tests on the circuits: RQT12.L8B1, RQS.L8B1, RCO.A78B1
DFBMA will me refilled in the afternoon
60A Closed Orbit Correctors: PCC and P2N whenever possible